OpenStudy (anonymous):

Could someone please explain the processes for these Algebra II function problems? 1. If f is a function such that f(x+2)=f(x)+f(2) show each of the following. a. f(0)=0 b. f(-2)=-f(2) 2. If f is a function such that f(-x)=-f(x), show that f(0)=0

OpenStudy (aum):

f(x+2) = f(x) + f(2) f(x) = f(x+2) - f(2) put x = 0 f(0) = f(2) - f(2) = 0 (proves a.) put x = -2 f(-2) = f(0) - f(2) = 0 - f(2) f(-2) = -f(2) (proves b.)

OpenStudy (aum):

2. f(-x) = -f(x) put x = 0 f(-0) = -f(0) -0 = 0. So f(0) = -f(0) f(0) = -1 * f(0) The only number that is unchanged when you multiply by -1 is zero. Therefore, f(0) = 0.